Direct material = $7
Direct Labor = $4
Variable Manufacturing Overhead = $2
Fixed Manufacturing Overhead = $3
Fixed Selling Expense = $3
Fixed Admin Expense = $2
Sales commission = $1
Variable Admin Expense $.50
(Costs are when 10,000 units are produced and sold)
Relevant range is 7,500-12,500 units

1) If 8,000 units are produced, what is the average fixed manufacturing cost per unit produced?

2) If 12,500 units are produced, what is the average fixed manufacturing cost per unit produced?

3) If 8,000 units are produced, what is the total amount of fixed manufacturing cost incurred to support this level of production?

4) If 12,500 units are produced, what is the total amount of fixed manufacturing cost incurred to support this level of production?

5) If 8,000 units are produced, what is the total amount of manufacturing overhead cost incurred to support this level of production? What is the total amount expressed in a per unit basis?

6) If 12,500 units are produced, what is the total amount of manufacturing overhead cost incurred to support this level of production? What is the total amount expressed in a per unit basis?

7) If the selling price is $22 per unit, what is the contribution margin per unit sold?

8) If 11,000 units are produced, what are the total amounts of direct and indirect manufacturing costs incurred to support this level of production?

9) What total incremental cost will be incurred if production increases from 10,000 to 10,001 units?
Answer #1
1
Fixed manufacturing cost 30000 =10000*3
Divide by Units produced 8000
Average fixed manufacturing cost per unit 3.75
2
Fixed manufacturing cost 30000 =10000*3
Divide by Units produced 12500
Average fixed manufacturing cost per unit 2.40
3
Total amount of fixed manufacturing cost 30000 =10000*3
4
Total amount of fixed manufacturing cost 30000 =10000*3
5
Variable manufacturing overhead 16000 =8000*2
Fixed manufacturing overhead 30000 =10000*3
Total amount of manufacturing overhead cost 46000
Manufacturing overhead per unit 5.75 =46000/8000
6
Variable manufacturing overhead 25000 =12500*2
Fixed manufacturing overhead 30000 =10000*3
Total amount of manufacturing overhead cost 55000
Manufacturing overhead per unit 4.40 =55000/12500
7
Selling price 22.00
Less: Variable costs
Direct materials 7.00
Direct labor 4.00
Variable manufacturing overhead 2.00
Sales commissions 1.00
Variable administrative expense 0.50
Total Variable costs 14.50
Contribution margin per unit sold 7.50
8
Direct materials 77000 =11000*7
Direct labor 44000 =11000*4
Total amount of direct manufacturing cost 121000
Variable manufacturing overhead 22000 =11000*2
Fixed manufacturing overhead 30000 =10000*3
Total amount of indirect manufacturing cost 52000
9
Direct materials 7.00
Direct labor 4.00
Variable manufacturing overhead 2.00
Incremental manufacturing cost 13.00